# Understanding Impact of Human Feedback via Influence Functions

### v1

@@ Abstract
In Reinforcement Learning from Human Feedback (RLHF), it is crucial to learn suitable reward models from human feedback to align large language models (LLMs) with human intentions. However, human feedback can often be noisy, inconsistent, or biased, especially when evaluating complex responses. Such feedback can lead to misaligned reward signals, potentially causing unintended side effects during the RLHF process. To address these challenges, we explore the use of influence functions to measure the impact of human feedback on the performance of reward models. We propose a compute-efficient approximation method that enables the application of influence functions to LLM-based reward models and large-scale preference datasets. In our experiments, we demonstrate two key applications of influence functions: (1) detecting common forms of labeler bias in human feedback datasets and (2) guiding labelers to refine their strategies to align more closely with expert feedback. By quantifying the impact of human feedback on reward models, we believe that influence functions can enhance feedback interpretability and contribute to scalable oversight in RLHF, helping labelers provide more accurate and consistent feedback. Source code is available at https://github.com/mintaywon/IF_RLHF.
@@ 1 Introduction
As large language models (LLMs) demonstrate remarkable capabilities across various domains, ensuring their behaviors align with human intentions becomes increasingly important. To this end, reinforcement learning from human feedback (RLHF) has emerged as a powerful solution for fine-tuning LLMs (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). In RLHF, human feedback is collected to train reward models that capture important human values, such as helpfulness and harmlessness (Bai et al., 2022a; Ji et al., 2024). LLMs are then fine-tuned to produce outputs that closely align with these reward models.
@@
In this work, we assess the impact of human feedback on reward models by utilizing influence functions (Hampel, 1974; Koh & Liang, 2017). However, a significant challenge arises when applying influence functions to reward models, especially large-parameter models like LLMs and those involving extensive preference datasets, due to the high computational costs involved. To address this, we introduce a compute-efficient method that utilizes vector compression techniques (Li & Li, 2023) alongside the influence estimation method (Kwon et al., 2024), achieving a 2.5-fold speed acceleration compared to previous methods in computing influence functions. This approach significantly reduces the computational costs required to compute influence functions, facilitating more practical applications in large-scale settings.
We demonstrate two applications of influence functions (see Figure 1 for an overview): (1) detecting labeler bias in training datasets, and (2) improving suboptimal labeling strategies. In our first experiment, we explore two prevalent biases in the RLHF paradigm: length and sycophancy bias, where labelers may naively prefer longer (Saito et al., 2023) and more sycophantic responses (Sharma et al., 2023), regardless of response quality. To test our approach, we introduce biased samples into the training dataset and assess whether they can be detected using influence functions. Our approach significantly outperforms several baselines, including GPT-4o (OpenAI, 2024) and various outlier detection methods (Lee et al., 2018), by effectively identifying biased samples.
@@ 2 Related work
Influence functions measure the impact of individual training data points on the resulting model and have been applied to various tasks, such as identifying influential data, detecting label errors, and interpreting model behavior (Koh & Liang, 2017; Guo et al., 2021; Kwon et al., 2024; Lin et al., 2024). Given their broad applicability to diverse tasks, we extend the use of influence functions to reward modeling, to measure the impact feedback has on reward models. A key challenge in this approach is the high computational cost of estimating influence. Building on recent advancements in efficient influence computation methods, which enables the estimation of influence functions even for LLMs (Kwon et al., 2024; Lin et al., 2024; Grosse et al., 2023), we apply influence functions to LLM-based reward models.
@@
As AI models become more powerful, reliably providing feedback on their behavior becomes increasingly challenging (Burns et al., 2024). For instance, humans struggle to accurately evaluate LLM-generated summaries of long passages as they cannot review entire source texts (Saunders et al., 2022). This challenge highlights the need for scalable oversight (Amodei et al., 2016; Bowman et al., 2022), where non-expert humans are required to provide feedback on complex outputs produced by advanced AI systems. A common approach to scalable oversight involves using capable AI models during the feedback process, either to assist humans (Saunders et al., 2022) or to replace them (Bai et al., 2022b; Cui et al., 2023). However, AI-assisted feedback processes can still fail, and it remains uncertain whether they will guarantee alignment (Hofstätter, 2023; Casper et al., 2023) for increasingly complex tasks. An alternative approach to scalable oversight is the “sandwich paradigm” (Cotra, 2021; Bowman et al., 2022), which places the capabilities of an LLM between a domain expert and the model overseer. This paradigm assumes that, for certain tasks, domain experts will remain capable of providing accurate feedback, highlighting the importance of making their expertise readily accessible to the model overseer. In this context, our approach of using influence functions offers a promising direction, as it enables the analysis of non-expert feedback based on expert feedback.
@@ 3.1 Influence functions
The influence function quantifies the impact of individual training data points on model parameters by measuring the change in parameters in response to an infinitesimal adjustment in the weight of a specific data point (Hampel, 1974; Koh & Liang, 2017). To be more specific, we denote a parameter by $\theta$, an associated parameter space by $\Theta$, a loss function by $\ell$, a parameterized model by $f_{\theta}$, and a training dataset by $\mathcal{D}$. The empirical risk minimizer $\theta^{*}$ is defined as $\theta^{*}:=\arg\min_{\theta\in\Theta}\ |\mathcal{D}|^{-1}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))$, and the $\varepsilon$-weighted risk minimizer for a single training data point $x_{i}\in\mathcal{D}$ is defined as follows:
$\theta^{(i)}(\varepsilon):=\underset{\theta\in\Theta}{\mathrm{arg\,min}}\ \frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))+\varepsilon\ell(f_{\theta}(x_{i})).$ (1)
The influence function is defined as the derivative of $\theta^{(i)}(\varepsilon)$ at $\varepsilon=0$, capturing how fast the parameter would change when the weight on $x_{i}$ is slightly changed. With the standard assumptions (e.g., twice-differentiability and strong convexity of a loss function $\ell$), the influence at training data point $x_{i}$ is expressed with the Hessian matrix of the empirical loss and the first-order gradient as follows (Cook & Weisberg, 1980):
$\mathcal{I}_{\theta^{*}}\left(x_{i}\right):=\left.\frac{d\theta^{(i)}(\varepsilon)}{d\varepsilon}\right|_{\varepsilon=0}=-H(\mathcal{D};\theta^{*})^{-1}\nabla_{\theta}\ell_{i}|_{\theta=\theta^{*}},$ (2)
where $H(\mathcal{D};\theta):=\nabla_{\theta}^{2}\left(\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))\right)$ and $\nabla_{\theta}\ell_{i}=\nabla_{\theta}\ell(f_{\theta}(x_{i}))$. In many recent machine learning applications, the focus has been extended beyond the model parameter to any univariate quantity of interest $f(\theta)$, such as validation loss or a model prediction, leading to the following influence function via the chain rule of derivatives (Koh & Liang, 2017):
@@
The influence function $\mathcal{I}_{f}(x_{i})$ quantifies the impact of a training data point $x_{i}$ on $f(\theta)$. Based on this derivation, it has been utilized in various downstream tasks such as detecting noisy labels (Koh & Liang, 2017; Pruthi et al., 2020; Guo et al., 2021) and interpreting model predictions (Han et al., 2020; Grosse et al., 2023).
@@ 3.2 Reinforcement learning from human feedback
RLHF is an effective technique for aligning LLMs with human preferences by incorporating human evaluations into the learning process. It has become increasingly standard due to its powerful capability to generate human-like, helpful, and safe model outcomes (Bai et al., 2022a; Ouyang et al., 2022; Dai et al., 2024). Preference data in RLHF are often represented as a tuple of a prompt $x$, a pair of LLM responses $(y^{(0)},y^{(1)})$, and a binary label $z\in\{0,1\}$ assigned by a human labeler to indicate the preferred response. For clarity, we introduce the notation $\mathbf{d}:=(x,y^{(0)},y^{(1)},z)$ to represent feedback data points. Such preference data are learned by minimizing the following cross-entropy loss based on the Bradley-Terry model (Bradley & Terry, 1952):
@@ 4.1 Influence functions in preference-based reward learning
$\mathcal{I}_{\tt val}(\mathbf{d}_{i}):=-\nabla_{\theta}\mathcal{L}(\mathcal{D}_{\tt val};\theta)^{\top}H_{\tt pref}(\mathcal{D}_{\tt tr};\theta)^{-1}\nabla_{\theta}\ell_{\tt pref}(\mathbf{d}_{i};\theta),$ (5)
@@
It is noteworthy that constructing targeted validation sets $\mathcal{D}_{\tt val}$ is crucial when utilizing influence functions, as they estimate the impact on validation loss. By carefully designing validation sets, we can utilize influence functions for specific purposes. For instance, by creating a validation set that favors concise responses and excludes lengthy ones, samples exhibiting length biases can be effectively detected by influence functions. Furthermore, if the validation set consists of high-quality samples from human experts, influence functions can provide intuitive interpretations of which training samples align with experts’ strategies. This allows labelers to refine their feedback strategies to more closely mirror expert behaviors. In our experiments, we demonstrate the diverse applications of influence functions based on the composition of the validation sets.
@@ 4.2 Efficient computation
$H_{\tt pref}(\mathcal{D}_{\tt tr};\theta)^{-1}$ $\approx-\frac{1}{n\lambda}\sum_{\mathbf{d}\in\mathcal{D}_{\tt tr}}\bigg{(}I-\frac{\nabla_{\theta}\ell_{\mathtt{pref}}(\mathbf{d};\theta)\nabla_{\theta}\ell_{\mathtt{pref}}(\mathbf{d};\theta)^{\top}}{\lambda+\nabla_{\theta}\ell_{\mathtt{pref}}(\mathbf{d};\theta)^{\top}\nabla_{\theta}\ell_{\mathtt{pref}}(\mathbf{d};\theta)}\bigg{)},$ (6)
where $\lambda>0$ is a small positive constant adopted during approximation. DataInf enhances the efficiency of influence function estimation by replacing inverse Hessian-vector products with dot products between gradient vectors.
However, DataInf requires significant storage capacity for large training datasets, as each gradient vector is as large as the model parameters $\theta$. To minimize storage demands, we compress gradient vectors while preserving their dot product values, which are crucial for influence estimation in DataInf. Inspired by Lin et al. (2024), we utilize the one-permutation one-random-projection (OPORP) method (Li & Li, 2023) to compress gradient vectors. Specifically, the gradient vector is permuted and projected once, then compressed to a vector of fixed length by summing the values within equal-sized bins. Using this procedure, we reduce the size of a single gradient vector from 160MB ($42M$ dimensions [^1] The gradient size is 42M due to the use of Low-Rank Adaptation (Hu et al., 2022) in reward modeling.) to 256KB ($65K$ dimensions), enabling the storage of entire gradients for large preference datasets. Influence estimation is significantly accelerated by utilizing this technique, as compression requires only one pass of backpropagation, and influence computation is completed within seconds using compressed gradients (see supporting results in Figure 5). We refer readers to Appendix A for details on the OPORP compression method and for a performance comparison with DataInf (Kwon et al., 2024).
@@ 5 Experiment
We design our experiments to investigate the following:
@@
5.1 Bias detection using influence functions
@@ 5.1.1 Experimental setup
We construct our training and validation sets using the helpful split of Anthropic’s Helpfulness-Harmlessness (Anthropic-HH) dataset (Bai et al., 2022a), which was annotated by humans who evaluated responses based on helpfulness, providing binary preference labels for conversations between a human and an assistant. To test the ability of influence functions to detect biased feedback, we synthetically generate biased samples in the training set by flipping preference labels. Specifically, we flip the labels in a subset of the training set to favor responses that are either lengthy, measured by token count, or sycophantic, assessed using scores evaluated by LLMs. [^2] Similar to the approach in  Sharma et al. (2023), we prompt LLMs to rate sycophancy and average these ratings to obtain a reference sycophancy score (see Appendix D for details). This manipulation affects 6.56% of the labels for the length bias experiments and 4.17% for the sycophancy bias experiments. Each training set comprises 15,000 samples.
@@
For both length and sycophancy bias experiments, we train reward models by fine-tuning the Llama-3-8B model (Dubey et al., 2024), appending a fully connected layer to the last token of the final layer. We utilize the trl library (von Werra et al., 2022) for reward model training. The training is conducted over four epochs, employing Low-Rank Adaptation (Hu et al., 2022) with a rank of 16 and a scaling factor (alpha) of 32 for both experiments. Training is conducted on a single NVIDIA RTX A6000 GPU.
@@
To detect biased samples using influence functions, we employ a threshold-based detector that classifies a training sample as biased if its influence score exceeds a specified threshold. We also consider baselines that utilize other metrics for scoring, such as Mahalanobis distance (Lee et al., 2018) and k-nearest neighbors (Sun et al., 2022), which measure the distance between a training sample and validation samples. Additionally, we use metrics like self-confidence and entropy to assess the prediction uncertainty of the reward model (Kuan & Mueller, 2022). Additionally, we evaluate LLM-based detectors, including GPT-4o (OpenAI, 2024) and Gemini-1.5-Pro (Reid et al., 2024), using few-shot prompting. Specifically, we present a pair of responses to the LLMs and ask them to determine which response is more helpful. Further details about these baselines are available in Appendix C.
@@
For evaluation, we compute the true positive rate (TPR) and false positive rate (FPR) using the threshold-based detector’s classification at different thresholds. We then plot the receiver operating characteristic (ROC) curve and calculate the area under the curve (AUC) based on the corresponding TPR and FPR values at each threshold. Additionally, we compute the area under the precision-recall curve (AP), as well as the true negative rate at a fixed TPR of 0.80 (TNR80). We report these metrics, along with the precision-recall curve in Appendix E.
Figure 2: ROC curves comparing influence detectors with baseline methods for detecting labeler biases: (left) length bias and (right) sycophancy bias. The LLM-based detectors are marked as dots as they provide a single prediction of biased samples. The dotted line represents performance at random (AUC = 0.5). Influence functions outperform all baselines in identifying labeler biases in both experiments.

### v2

@@ Abstract
In Reinforcement Learning from Human Feedback (RLHF), it is crucial to learn suitable reward models from human feedback to align large language models (LLMs) with human intentions. However, human feedback can often be noisy, inconsistent, or biased, especially when evaluating complex responses. Such feedback can lead to misaligned reward signals, potentially causing unintended side effects during the RLHF process. To address these challenges, we explore the use of influence functions to measure the impact of human feedback on the performance of reward models. We propose a compute-efficient approximation method that enables the application of influence functions to LLM-based reward models and large-scale preference datasets. Our experiments showcase two key applications of influence functions: (1) detecting common labeler biases in human feedback datasets and (2) guiding labelers in refining their strategies to better align with expert feedback. By quantifying the impact of human feedback, we believe that influence functions can enhance feedback interpretability and contribute to scalable oversight in RLHF, helping labelers provide more accurate and consistent feedback. Source code is available at https://github.com/mintaywon/IF_RLHF.
Understanding Impact of Human Feedback via Influence Functions
Taywon Min1, Haeone Lee1, Yongchan Kwon2, Kimin Lee1 1KAIST; 2Columbia University kiminlee@kaist.ac.kr
@@ 1 Introduction
As large language models (LLMs) demonstrate remarkable capabilities across various domains, ensuring their behaviors align with human intentions becomes increasingly important. To this end, reinforcement learning from human feedback (RLHF) has emerged as a powerful solution for fine-tuning LLMs (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). In RLHF, human feedback is collected to train reward models that capture important human values, such as helpfulness and harmlessness (Bai et al., 2022; Ji et al., 2024). LLMs are then fine-tuned to produce outputs that closely align with these reward models.
@@
In this work, we assess the impact of human feedback on reward models by utilizing influence functions (Hampel, 1974; Koh and Liang, 2017). However, a significant challenge arises when applying influence functions to reward models, especially large-parameter models like LLMs and those involving extensive preference datasets, due to the high computational costs involved. To address this, we introduce a compute-efficient method that utilizes vector compression techniques (Li and Li, 2023) alongside the influence estimation method (Kwon et al., 2024), achieving a 2.5-fold speed acceleration compared to previous methods in computing influence functions. This approach significantly reduces the computational costs required to compute influence functions, facilitating more practical applications in large-scale settings.
We demonstrate two applications of influence functions (see Figure 1 for an overview): (1) detecting labeler bias in training datasets, and (2) improving suboptimal labeling strategies. In our first experiment, we explore two prevalent biases in the RLHF paradigm: length and sycophancy bias, where labelers may naively prefer longer (Saito et al., 2023) and more sycophantic responses (Sharma et al., 2023), regardless of response quality. To evaluate our approach, we introduce labeler bias by flipping a small portion of the Anthropic-HH dataset Bai et al. (2022) and assess whether they can be detected using influence functions. Our approach significantly outperforms several baselines, including GPT-4o (OpenAI, 2024) and various outlier detection methods (Lee et al., 2018; Sun et al., 2022), by effectively identifying biased samples. We further extend our analysis to original Anthropic-HH datasets (i.e., without flipping) using human surveys, confirming its robustness. Specifically, when analyzing 100 samples flagged by our approach, human annotators found 47 mislabeled cases, indicating the prevalence of labeler bias in real-world datasets. These results demonstrate the potential of our method to enhance publicly available preference datasets by accurately detecting and mitigating labeler biases.
@@ 2 Related Work
Influence functions originate from robust statistics (Hampel, 1974), which aim to robustly fit models based on the majority of the data, allowing accurate detection of outliers (Tyler, 2008; Rousseeuw and Leroy, 2003). Leveraging this property, influence functions have been applied to Deep Neural Networks to estimate the impact of data points on model behavior (Koh and Liang, 2017). Specifically, it has been used on image datasets to detect adversarial samples or augmenting data (Lee et al., 2020; Cohen et al., 2020). It has also been applied to text classification tasks (Schioppa et al., 2022; Guo et al., 2021). Given their broad applicability, we extend the use of influence functions to reward modeling, specifically to measure the impact of human feedback on reward models. We build on recent advancements in efficient influence computation methods, which enable the estimation of influence functions even for large language models (LLMs) (Kwon et al., 2024; Lin et al., 2024; Grosse et al., 2023). Based on DataInf (Kwon et al., 2024), we apply influence functions to LLM-based reward models to analyze labeler bias and improve alignment in reward modeling.
@@
As AI models become more powerful, reliably providing feedback on their behavior becomes increasingly challenging (Burns et al., 2024). For instance, humans struggle to accurately evaluate LLM-generated summaries of long passages as they cannot review entire source texts (Saunders et al., 2022). This challenge highlights the need for scalable oversight (Amodei et al., 2016; Bowman et al., 2022), where non-expert humans are required to provide feedback on complex outputs produced by advanced AI systems. A common approach to scalable oversight involves using capable AI models during the feedback process, either to assist humans (Saunders et al., 2022) or to replace them (Bai et al., 2022; Cui et al., 2023). However, AI-assisted feedback processes can still fail, and it remains uncertain whether they will guarantee alignment (Hofstätter, 2023; Casper et al., 2023) for increasingly complex tasks. An alternative approach to scalable oversight is the “sandwich paradigm" (Cotra, 2021; Bowman et al., 2022), which places the capabilities of an LLM between a domain expert and the model overseer. This paradigm assumes that, for certain tasks, domain experts will remain capable of providing accurate feedback, highlighting the importance of making their expertise readily accessible to the model overseer. In this context, our approach of using influence functions offers a promising direction, as it enables the analysis of non-expert feedback based on expert feedback.
@@ 3.1 Influence Functions
The influence function quantifies the impact of individual training data points on model parameters by measuring the change in parameters in response to an infinitesimal adjustment in the weight of a specific data point (Hampel, 1974; Koh and Liang, 2017). To be more specific, we denote a parameter by $\theta$, an associated parameter space by $\Theta$, a loss function by $\ell$, a parameterized model by $f_{\theta}$, and a training dataset by $\mathcal{D}$. The empirical risk minimizer $\theta^{*}$ is defined as $\theta^{*}:=\arg\min_{\theta\in\Theta}\ |\mathcal{D}|^{-1}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))$, and the $\varepsilon$-weighted risk minimizer for a single training data point $x_{i}\in\mathcal{D}$ is defined as follows:
$\theta^{(i)}(\varepsilon):=\underset{\theta\in\Theta}{\mathrm{arg\,min}}\ \frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))+\varepsilon\ell(f_{\theta}(x_{i})).$
The influence function is defined as the derivative of $\theta^{(i)}(\varepsilon)$ at $\varepsilon=0$, capturing how fast the parameter would change when the weight on $x_{i}$ is slightly changed. With the standard assumptions (e.g., twice-differentiability and strong convexity of a loss function $\ell$), the influence at training data point $x_{i}$ is expressed with the Hessian matrix of the empirical loss and the first-order gradient as follows (Cook and Weisberg, 1980):
$\mathcal{I}_{\theta^{*}}\left(x_{i}\right):=$ $\left.\frac{d\theta^{(i)}(\varepsilon)}{d\varepsilon}\right|_{\varepsilon=0}$
$=$ $-H(\mathcal{D};\theta^{*})^{-1}\nabla_{\theta}\ell_{i}|_{\theta=\theta^{*}},$ (1)
where $H(\mathcal{D};\theta):=\nabla_{\theta}^{2}\left(\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\ell(f_{\theta}(x))\right)$ and $\nabla_{\theta}\ell_{i}=\nabla_{\theta}\ell(f_{\theta}(x_{i}))$. In many recent machine learning applications, the focus has been extended beyond the model parameter to any univariate quantity of interest $f(\theta)$, such as validation loss or a model prediction, leading to the following influence function via the chain rule of derivatives (Koh and Liang, 2017):
@@
The influence function $\mathcal{I}_{f}(x_{i})$ quantifies the impact of a training data point $x_{i}$ on $f(\theta)$. Based on this derivation, it has been utilized in various downstream tasks such as detecting noisy labels (Koh and Liang, 2017; Pruthi et al., 2020; Guo et al., 2021) and interpreting model predictions (Han et al., 2020; Grosse et al., 2023).
@@ 3.2 Reinforcement Learning from Human Feedback
RLHF is an effective technique for aligning LLMs with human preferences by incorporating human evaluations into the learning process. It has become increasingly standard due to its powerful capability to generate human-like, helpful, and safe model outcomes (Bai et al., 2022; Ouyang et al., 2022; Dai et al., 2024). Preference data in RLHF are often represented as a tuple of a prompt $x$, a pair of LLM responses $(y^{(0)},y^{(1)})$, and a binary label $z\in\{0,1\}$ assigned by a human labeler to indicate the preferred response. For clarity, we introduce the notation $\mathbf{d}:=(x,y^{(0)},y^{(1)},z)$ to represent feedback data points. Such preference data are learned by minimizing the following cross-entropy loss based on the Bradley-Terry model (Bradley and Terry, 1952):
@@ 4.1 Influence Functions in Preference-Based Reward Learning
$\mathcal{I}_{\tt val}(\mathbf{d}_{i}):=$ $-\nabla_{\theta}\mathcal{L}(\mathcal{D}_{\tt val};\theta)^{\top}$
$H_{\tt pref}(\mathcal{D}_{\tt tr};\theta)^{-1}\nabla_{\theta}\ell_{\tt pref}(\mathbf{d}_{i};\theta),$
@@
Constructing targeted validation sets $\mathcal{D}_{\tt val}$ is crucial when utilizing influence functions, as they estimate the impact on validation loss. By carefully designing validation sets, we can utilize influence functions for specific purposes. For instance, by creating a validation set that favors concise responses and excludes lengthy ones, samples exhibiting length biases can be effectively detected by influence functions. Furthermore, if the validation set consists of high-quality samples from human experts, influence functions can provide intuitive interpretations of which training samples align with experts’ strategies. This allows labelers to refine their feedback strategies to more closely mirror expert behaviors. In our experiments, we demonstrate diverse applications of influence functions based on the composition of validation sets.
@@ 4.2 Efficient Computation
$H_{\tt pref}(\mathcal{D}_{\tt tr};\theta)^{-1}$ $\approx\frac{1}{n\lambda}\sum_{\mathbf{d}\in\mathcal{D}_{\tt tr}}\left(I-\frac{v_{\mathbf{d}}v_{\mathbf{d}}^{\top}}{\lambda+v_{\mathbf{d}}^{\top}v_{\mathbf{d}}}\right),$
where $\lambda>0$ is a positive constant adopted during approximation and $v_{\mathbf{d}}=\nabla_{\theta}\ell_{\mathtt{pref}}(\mathbf{d};\theta)$. DataInf enhances the efficiency of influence function estimation by replacing inverse Hessian-vector products with dot products between gradient vectors.
However, DataInf requires significant storage capacity for large training datasets, as each gradient vector is as large as the model parameters $\theta$. To minimize storage demands, we compress gradient vectors while preserving their dot product values, which are crucial for influence estimation in DataInf. Inspired by Lin et al. (2024), we utilize the one-permutation one-random-projection (OPORP) method (Li and Li, 2023) to compress gradient vectors. Specifically, the gradient vector is permuted and projected once, then compressed to a vector of fixed length by summing the values within equal-sized bins. Using this procedure, we reduce the size of a single gradient vector from 160MB ($42M$ dimensions [^1] The gradient size is 42M due to the use of Low-Rank Adaptation (Hu et al., 2022) in reward modeling.) to 256KB ($65K$ dimensions), enabling the storage of entire gradients for large preference datasets. Influence estimation is significantly accelerated by utilizing this technique, as compression requires only one pass of backpropagation, and influence computation is completed within seconds using compressed gradients (see supporting results in Figure 3). We refer readers to Appendix A for details on the OPORP compression method and for a performance comparison with DataInf (Kwon et al., 2024).
@@ 5 Experiment
We design our experiments to investigate the following questions:
@@
5.1 Bias Detection
@@ 5.1.1 Experimental Setup
We construct our training and validation sets using the helpful split of Anthropic’s Helpfulness-Harmlessness (Anthropic-HH) dataset (Bai et al., 2022), where human annotators provided binary preference labels based on response helpfulness in human-assistant conversations. To assess the ability of influence functions to detect biased feedback, we synthetically generate biased samples in the training set by flipping preference labels. Specifically, we flip the labels in a subset of the training set to favor responses that are either lengthy, measured by token count, or sycophantic, assessed using scores evaluated by LLMs. [^2] Similar to the approach in  Sharma et al. (2023), we prompt LLMs to rate sycophancy and average these ratings to obtain a reference sycophancy score (see Appendix D for details). This manipulation affects 6.56% of the labels for the length bias experiments and 4.17% for the sycophancy bias experiments. Each training set comprises 15,000 samples.
@@
For both length and sycophancy bias experiments, we train reward models by fine-tuning the Llama-3-8B model (Dubey et al., 2024), appending a fully connected layer to the last token of the final layer. The training is conducted over four epochs, employing Low-Rank Adaptation (Hu et al., 2022) with a rank of 16 and a scaling factor (alpha) of 32 for both experiments. Training is conducted on a single NVIDIA RTX A6000 GPU.
Figure 2: ROC curves comparing influence detectors with baseline methods for detecting (left) length bias and (right) sycophancy bias. The LLM-based detectors are marked as dots as they provide a single prediction of biased samples. The gray dotted line represents performance at random (AUC = 0.5). Influence functions outperform all baselines in both experiments.
@@
To detect biased samples using influence functions, we employ a threshold-based detector that classifies a training sample as biased if its influence score exceeds a specified threshold. We also consider baselines that utilize other metrics for scoring, such as Mahalanobis distance (Lee et al., 2018; Bai et al., 2022) and k-nearest neighbors (Sun et al., 2022), which measure the distance between a training sample and validation samples. Additionally, we use self-confidence and entropy metrics to assess the prediction uncertainty of the reward model (Kuan and Mueller, 2022). Finally, we evaluate LLM-based detectors, including GPT-4o (OpenAI, 2024) and Gemini-1.5-Pro (Reid et al., 2024), using few-shot prompting. Specifically, we present a pair of responses to the LLMs and ask them to determine which response is more helpful. Further details about these baselines are available in Appendix C.
@@
For evaluation, we compute the true positive rate (TPR) and false positive rate (FPR) using the threshold-based detector’s classification at different thresholds. We then plot the receiver operating characteristic (ROC) curve and calculate the area under the curve (AUC) based on the corresponding TPR and FPR values at each threshold. Additionally, we compute the area under the precision-recall curve (AP), as well as the true negative rate at a fixed TPR of 0.80 (TNR80). These metrics, along with the precision-recall curve, are reported in Appendix E.